Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

19q16

Expert replies
by stubbornp » Sat Oct 18, 2008 10:32 pm
16. What is the probability that events A and B both occur?
(1) The probability that event A occurs is 0.8.
(2) The probability that event B occurs is 0.6.
Join the discussion
Source: — Data Sufficiency |

by Saffa » Sun Oct 19, 2008 12:12 am
Funny one: Surely the option are only C and E;

I would say 0.6 x 0.8 = 0.42 thus SUFF

IMO C
Join the discussion

by stubbornp » Sun Oct 19, 2008 5:23 pm
its wrong...Any other volunteer??
Join the discussion

by Ian Stewart » Sun Oct 19, 2008 6:50 pm
The answer is E; the probability that *both* occur is only equal to 0.6*0.8 if the two events are independent (and Saffa- 0.6*0.8 is 0.48, not 0.42 :) ). Imagine the following problem:

Pick a random integer between 1 and 5 inclusive.
Let A be "the number is less than or equal to 4"
Let B be "the number is less than or equal to 3"

Clearly the probability that both A and B occur is 0.6, not 0.48.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by stubbornp » Sun Oct 19, 2008 10:16 pm
Ian Stewart wrote:The answer is E; the probability that *both* occur is only equal to 0.6*0.8 if the two events are independent (and Saffa- 0.6*0.8 is 0.48, not 0.42 :) ). Imagine the following problem:

Pick a random integer between 1 and 5 inclusive.
Let A be "the number is less than or equal to 4"
Let B be "the number is less than or equal to 3"

Clearly the probability that both A and B occur is 0.6, not 0.48.
Awsome explanation....tnx ian...
Join the discussion